May June 2012

From NPDGamma Wiki

Jump to: navigation, search

Contents

May -16-2013 - Plans for completion of run and tasks during maintenance break

  • With beam:
    • Polarimetry with empty target
  • Without beam:
    • Measure noise asymmetry-DONE
    • From data of noise measurements determine RMS noise width per detector and compare with results from earlier measurements. SEEMS TO BE SIGNIFICANTLY SMALLER RMS=30-50 MIROVOLTS
    • Calibrate M1 and M4 preamplifiers; stability and linearity. DONE ELOG HAS RESULTS
    • Replace Clover. DONE
    • Check a few 45V batteries in detector and also M1 and M4 HV batteries. NOT YET
    • Do we have plans for VMEs??
    • Schedule for end of run (pdf)

June -16-2012 - A start of a game plan for target testing

  1. This is a only a few ideas what we could do with the target
    1. The LH2 target provides the experiment para hydrogen. The experiments needs to know from us the hydrogen number density and the ortho/para ratio. For the number density we need the density of the target defined by temperature of the target. The o/p ratio we should get from neutron transmission of target in/out and normalized with M2. We need to think how we define the target temperature which then would give the density. It turned out that temperature dependence of the density at these low temperatures is not very strong.

What is the target temperature? If we look at different sensors we see that we have gradients if we believe the sensors are accurate - so what is temperature of the target, we need to give a single value to experiment. One way to test the gradients is to make small change(s) to target temperature and see how the different sensors are responding.

    1. We could run opc lower or higher temperatures and see what the target will end up - to another stable condition
    2. we have already operated the system when xxx refrigerator was off so we know that mode
    3. what about switching precooler refrigerator off - during the next run that would save this compressor if we do not need to use it all the time.
    4.  ????

June -06-2012 - Addition to run/game plan

  1. beam with low power, hydrogen target density measurements
    1. beam at low power (350 kW). Not worth taking production LH2 data,
    2. worth to change density around and see the behavior of target.
  2. Additional tasks with beam
    1. Testing of Bessel filter/amplifier circuit
    2. Move M4 next to cryostat and see how much we gain spectrum by shortening the flight path length. We need more data points to improve statistics in OP ratio measurement.
  3. Additions for game plan without beam are:
    1. To check linearity of the gamma detector front electronics using LEds and one channel. This still needs to be thought through how to do this and what would sensitivity to be.
    2. For false asymmetry study from M1, use M1 with M4 preamplifier. For charge injection to the preamp use battery and resistor. Then with LEDs on check multiplicative false asymmetry and LEDs off the additive asymmetry. The rest of the DAQ should be running as was operated during runs with beam.

May-31-2012 - A Run Plan till Jun-18 and over

This modified list is based on the circulated list by Mike and discussed by EC.

  1. May 30-June 1: 3 days of beam on LH2. A good decay run at the end.
  2. June 2: A day of beam on Cl ( gets data for paper, brackets LH2 data on both ends, confirms polarization,..) we need most probably collimate beam more to reduce rad doses – can we reach any interesting statistics?
    • decay run
  3. June 3-5: 3 days of accelerator work/tuning
    1. more noise runs ?
    2. more LED runs ?
    3. OR if the Bessel filter circuit is available make runs with Bessel filter – noise runs and LED runs and be ready with filter when beam comes to take data for a day?
  4. June 6-11: 6 days of beam:
    1. start Al measurement – we will have very low count rate since the new Al tgt is thin 14% from old one. First step is install the Al tgt without back Li absorber and with 3He see that we have known flipping ratio. This check is to see that Al tgt doesn’t disturb the RF field and thus reduce the flipping ratio (Al target is forced to be a non-stopping target to minimize spin flipper perturbation so enough beam gets through to do 3He).
    2. If flipping ratio is fine we need to install the Li shield to prevent scattered and transmitted neutrons to get into LH2. Then we can run 2-3 days with Al.
    3. A good Al decay run
    4. The mount CsI tgt and run long enough to get bound for asymmetry, this can take 2-3 days
    5. A good CsI decay run
    6. If time left run LH2
  5. June 12: a scheduled maintenance day
  6. June 13-17: 5 days of beam: how to use this time depends on results from first 2 weeks:
    1. a. Bessel filters are highest priority. We will have electronics for one ring, hard to get 10e-8 level measurement however we need to run enough, maybe total of 2 days with M1 connected
    2. We can take more data on LH2 – 3 days left
    3. More electronic asymmetry measurement with beam absorbed and M1 connected – to see that electronic asymmetry is gone
    4. More electronic asymmetry data with beam absorbed and M1 connected and LEDs on – to see that electronic asymmetry is gone
    5. Calibrating M4 for tgt density changes. Warm tgt up controlled way and see effect on transmitted flux. 2-4 h
  7. After beam turned off
    1. Do we need to measure more noise or LEDs ?

Draft NPDG game plan for the May/15 and after

May/15 ->May/22 beam at .9 MW. At present we have 3 days of hydrogen production. Kyle reports rough raw asymmetry

Aud = -4.985e-8 +- 9.51e-8

Alr = -7.095e-8 +- 9.48e-8

offset = -9.826e-9 +- 5.79e-8 .

The near-term/longer-term (till Mid June) strategy is to make no changes to the present apparatus and double the amount of LH data in order to obtain a PV asymmetry with an uncertainty of ~ 7 e-8. In addition we need to perform systematics checks to consolidate the result before the mid June.

A polarized 3He cell will be available on 5/16 and we plan to measure the flipping ration in order to establish that the beam is polarized and the spin flipper is working properly. After He-3 study we will measure the Cl asymmetry and then resume the hydrogen production till mid June. We nee to discuss if during next month or so we need to do more He-3 studies and Cl runs.

In the next few days the following systematic checks we need to perform are:

  1. 1. Close shutter during steady beam (28Al ground state (2 min half life) saturated) to determine the beta delayed and prompt signal level from Al. The ratio of beta-delayed gammas to prompt gammas is 1.79/8.75. We expect that the ratio of prompt Al to prompt capture on LH will be ~ 15%. DONE 5/14 and again 5/15
  2. 2. When beam goes off continue taking data to measure for long-lived beta delayed gammas.
  3. 3. Beam off additive asymmetry measurement for 6 hours
  4. 4. Beam off LED on multiplicative asymmetry measurement for 6 hours.
  5. 5. Repeat 3. and 4. With beam on and 2 layers of Li blocking beam. There is a large (1 e-5) asymmetry in the monitor signal with beam on. We need to establish that this large spin-flipper-state signal is not somehow producing a false asymmetry.
  6. 6. Obtain a total of 6 days LH data
  7. 7. 6 hours Cl in spin flipper to establish that we can measure a PV asymmetry.

The above should give an uncertainty in the PV asymmetry of ~ 7 e-8.

Prepare for a second cycle of data and systematics to reduce the uncertainty to 5 e-8. The only change would be to power the monitors, or at least one monitor, with batteries in order to break the ground loop formed by the signal cable and the shield of the HV cable.

Discussions with Scott Wilburn indicate that we may have 4 Bessel-filter-amplifiers in 2 weeks. We would then test the BFA and if they work take data using the BFA’s.

Personal tools